# Logic Synthesis Optimization with Predictive Self-Supervision via Causal Transformers

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
logic synthesis optimization (LSO) in which a logic circuit is transformed into an functionally equivalent and optimized representation is a part of EDA workflow (Sasao, 1993). In LSO, a sequence of optimization transformations supported by academia and industry (Amarú et al., 2017) and named recipe are applied to logic-level circuit designs to optimize performance criteria such as size and speed, named Quality of Results (QoR), in the final IC.
@@
Due to high latency and long running time of EDA tools, there is an huge surge of interest to use Machine Learning (ML) for EDA (Yang et al., 2022c; Huang et al., 2021; Ghose et al., 2021; Chowdhury et al., 2023b; Sohrabizadeh et al., 2023), specifically ML-guided logic synthesis (Huang et al., 2021), (Yang et al., 2022b). In addressing this, logic synthesis designers frequently employ policy-based Reinforcement Learning (RL) methods to identify optimal sequences of optimizations (Grosnit et al., 2022). Given the time-intensive nature of these searches, leveraging pre-trained policy functions can expedite RL processes (Chowdhury et al., 2024). This has led researchers to explore new challenges in developing ML models that exhibit robust generalization capabilities for predicting the final QoR based on gate-level circuit representations and optimization sequences (Chowdhury et al., 2021).
@@
Most models in this domain input logic-level circuit designs in And-Inverter Graph (AIG) format (Mishchenko et al., 2006), capitalizing on the inductive bias inherent in the Directed Acyclic Graph (DAG) structure (Neto et al., 2019) whose nodes are partially ordered. Significant research is dedicated to inject the inductive bias of DAGs into message-passing graph encoders (Thost and Chen, 2021) or Graph Transformers (Luo et al., 2024) to enhance expressiveness. One contribution of our paper builds on top of these concepts, effectively generating graph-level representations that are well-suited for AIGs.
@@
The current LSO models often experience overfitting due to a limited number of publicly available circuits used during training, resulting in a lack of generalization, especially when tested on unseen circuits that differ in size and characteristics from the training circuits. Additionally, these models struggle with imbalanced input samples across different modalities; typically, there are more recipe available than AIGs. This disparity can lead to prominent overfitting, particularly due to the low expressiveness of graph encoders, which becomes exacerbated when the model excessively adapts to the encoder part handling recipe.
@@
Recently, self-supervised learning has addressed both the scarcity of data, offering an alternative to costly data augmentation methods, and issues of expressiveness in various downstream tasks related to LSO (Chowdhury et al., 2023a; Wang et al., 2022). This approach involves pre-training the circuit encoder component in advance, followed by fine-tuning it on specific downstream tasks. Inspired by predictive SSL (Liu et al., 2022), another contribution of our work is the definition of an auxiliary task that involves predicting intermediate QoRs. This joint training mechanism provides a stronger supervisory signal during training, enhancing the learning process.
@@
Another drawback of existing method is that the encoder used for encoding recipe lacks information from the input graphs, leading to an information bottleneck. This bottleneck occurs particularly when graph embeddings and recipe embeddings are merged using standard concatenation techniques. This limitation motivated us to develop a decoder-only transformer model for training, employing a joint loss function. As a result, the model contextualizes the embeddings of heuristics using AIG embeddings through a Cross-Attention module, facilitating more effective integration of information.
@@ 2. Related Work
Several policy-based RL methodologies, such as those discussed in (Yang et al., 2022b; Chowdhury et al., 2024; Wu et al., 2022b), leverage a pre-trained function that correlates initial circuits and recipe with the ultimate QoR. This architecture is split into two primary branches: one for encoding AIGs and another for recipes, along with a fusion branch that maps the embeddings onto the predicted QoR. Similar efforts to forecast QoRs based on initial graphs and recipe are documented in (Chowdhury et al., 2021; Yang et al., 2022b, a; Wu et al., 2022a, b; Zheng et al., 2024), where various graph and sequence encoders are explored. Discussions abound concerning the most effective AIG encoders and methods for integrating sequence embeddings with graph embeddings. As a result, improvements in encoder and fusion modules are crucial for enhancing prediction performance. For example, (Chowdhury et al., 2021) utilizes a GCN to encode graphs and a CNN for recipes, employing both mean and max pooling for graph combination. Meanwhile, (Yang et al., 2022a) employs GraphSage for graph encoding and a transformer for recipe encoding. Similar to one of recent baselines, LOSTIN (Wu et al., 2022a), (Yang et al., 2022b) integrates a Graph Isomorphism Network (GINE) for graph learning with an LSTM for recipe encoding. Additionally, (Wu et al., 2022b) explores various graph encoders, such as PNA and pooling techniques, alongside virtual supernodes in conjunction with hierarchical learning and RL methods. In all aforementioned methods, recipe embeddings and Graph embeddings are fused using concatenation. The MLP decoder has been used to generate the predicted QoR after fusion. However, none of these models utilize an attention mechanism to fuse the two branches, which would contextualize the embeddings for more accurate final predictions. There are several attempts toward fast and effiecient LSO models
@@
$y^{k}$ The ground truth QoR of $k^{th}$ step
@@ 3.2.2. Final QoR prediction
(2) $loss=\sum_{k=1}^{M}MSE(\hat{y}^{k},y^{k})$
Where $y^{k}$ represents the ground truth QoR of $k^{th}$ step.
Figure 1. Schematic depiction of the model architecture.
@@ 3.3.2. Level-wise Node Pooling
Table 2. Performance comparison of prediction models for delay and area across datasets. The models are evaluated using MAPE (%) and examined under two recipe and IP inductive setup.
@@ 3.3.4. Transformer Decoder
The mask $\bar{M}$ is typically an upper triangular matrix with zeros on and below the diagonal and $\infty$- ∞ above the diagonal for each sequence in the batch. This structure effectively impose causality and ensures that each position in the sequence can only attend to itself and previous positions, not to any future positions.
@@
Where $\widetilde{h}^{i}=(\widetilde{u}_{1},\widetilde{u}_{2},...,\widetilde{u}_{M})$ is the output sequence of embeddings for input AIG and $i^{th}$ recipe; $\widetilde{u}_{i}\in\mathbb{R}^{2\times d_{h}}$ represents generated embedding correspoding to $i^{th}$ QoR. Add & Norm: Each sub-layer (self-attention and feed-forward network) includes a residual connection followed by layer normalization:
@@
(14) $\hat{y}^{i}=MLP(\widetilde{h}^{i})$
Table 3. Statistics for Benchmark Datasets
@@ 4.2. Datasets
Three datasets have been employed for the experiments, each comprising a collection of circuit designs along with a set of recipe. Since each dataset is utilized in different studies, we have opted not to combine datasets in order to simplify comparisons with counterpart baselines. Additionally, we report performance metrics individually for each dataset. A consistent set of 1500 unique recipes borrowed from (Chowdhury et al., 2021) was applied across all datasets. Each recipe was constructed utilizing seven primary heuristics and their associated flags from the ABC toolkit. These heuristics include Balance, Rewrite (rw, rw -z), Refactor (rf, rf -z), and Re-substitution (rs, rs -z).
@@
Table 4. Ablation Study on Recipe Enc. and AIG Enc. using Freeze Probing setup. (Enc.: Encoder)
@@
Table 5. Ablation Study on decoder and SSL auxilary task. P.Q.T. stands for Predicting QoR trajectory.
@@ 4.3. Experimental Protocol
The efficacy of our model is assessed under both transductive and inductive settings to demonstrate its superiority over competing architectures. The dataset is split into training and validation sets with ratios of 0.66 and 0.33, respectively, for both setups. Additionally, zero-mean normalization has been applied to normalize the final QORs. Additionally, intermediate QoRs are normalized using the mean and standard deviation derived from the final QoRs. In the ablation study, we focus exclusively on the inductive setting, which presents the most challenging scenario for the LSO problem. In these experiments, given the challenges associated with predicting timing, delay has been selected as the target variable for QoR prediction.
Inductive Setup (Circuit-wise Inductive): In this setup, the circuits used during the training phase are different from those employed in testing. It is noteworthy that all recipes from the sample set are observed and applied to the circuits introduced to the model during training.
Transductive Setup (Recipe-wise Inductive): In the transductive setup, each circuit is exposed during training, but separate recipes are used for testing. Specifically, the set of recipes is partitioned into test and validation subsets, and all Circuits are visible during both training and testing phases, albeit with different recipes.
@@ 4.4. Results
Table 6. Ablation Study on different decoder layer for auto-regressive and causal QoR trajectory prediction
@@ 4.4.1. Comparison Study Result
In Table 2, we evaluate the MAPE performance of various QOR prediction models against the baseline in multi trials across two setups for predicting delay and area. The baseline model is the architecture proposed in the OABCD paper (Chowdhury et al., 2021). In the IP-inductive setup, the LSOformer surpasses all existing models and outperforms the baseline by 5.74%, 4.35%, and 17.06% for the EPFL, OABCD, and PD datasets, respectively, in terms of delay prediction. For area prediction, LSOformer reduces the MAPE across all datasets compared to the baseline and nearly all models. However, on the OABCD dataset, the GNN-H model achieves the highest performance, demonstrating a 14% improvement over the baseline. This pattern holds in the recipe-inductive setup for both delay and area predictions across the three datasets, where LSOformer exceeds the baseline by 11.40%, 2.38%, and 12.82% for the EPFL, OABCD, and PD datasets in delay prediction, respectively.
@@ 4.4.2. Ablation Study
Figure 2. Mean and circuit-wise performance comparison across test Circuits on Area (a) and Delay (b)
We assess the impact of the SSL auxilary loss integratd to the training pipeline. Similar to the benchmark SSL works (Liu et al., 2022), for the random setup, different parts of architecture are freezed to generate the lower bound of performance for SSL task. The difference between the results of random encoders and supervised models is considered the gap that the frozen model can potentially bridge in the best-case scenarios. Freezed LSOformer in Table 4, benefiting from QoR trajectory prediction, has performance near that of the supervised LSOformer and shows better results compared to the baseline, although it has not trained on final QoR.
@@
The effects of varying the embedding fusion and decoder configurations using different sequence decoders are explored in this work. The baseline model, employing a shared MLP in conjunction with concatenated recipe and AIG embeddings for predicting all intermediate and final QORs, underperforms relative to alternatives. Separate MLPs for each QoR from initial to final are also deployed, mimicking a multi-task setup. This configuration significantly enhances performance across all datasets compared to the baseline. Additionally, we experimented with an auto-regressive LSTM module designed to train and infer models auto-regressively, thereby introducing causality to LSTM. This model demonstrates improvements over the baseline in all datasets except for OABCD. As can be seen in Table 6, the LSOformer surpasses both the baseline and all other methods across the three datasets, affirming the efficacy of our architectural design choices.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: The Design Automation Conference; June 22–25, 2025; San Francisco, CAisbn: 978-1-4503-XXXX-X/18/06ccs: Hardware Electronic design automation (EDA); Logic synthesisccs: Computing methodologies Machine learning Machine learning approaches
@@ 1. Introduction
logic synthesis optimization (LSO) in which a logic circuit is transformed into a functionally equivalent and optimized representation is a part of EDA workflow (Sasao, 1993). In LSO, a sequence of optimization transformations supported by academia and industry (Amarú et al., 2017) and named recipe are applied to logic-level circuit designs to optimize performance criteria such as size and speed, named Quality of Results (QoR), in the final IC.
@@
Due to high latency and long running time of EDA tools, there is a huge surge of interest to use Machine Learning (ML) for EDA (Yang et al., 2022c; Huang et al., 2021; Ghose et al., 2021; Chowdhury et al., 2023b; Sohrabizadeh et al., 2023), specifically ML-guided logic synthesis (Huang et al., 2021), (Yang et al., 2022b). In addressing this, logic synthesis designers frequently employ policy-based Reinforcement Learning (RL) methods to identify optimal recipes (Grosnit et al., 2022). Given the time-intensive nature of searches, leveraging pre-trained policy functions can expedite RL processes (Chowdhury et al., 2024). This has introduced a new research challenge: developing ML models that act as policy functions or serve as alternatives to running costly EDA pipelines, capable of accurately mapping a pair of gate-level circuit representations and a recipe to the final QoR (Chowdhury et al., 2021).
@@
Most models in this domain represent logic-level circuit designs using the And-Inverter Graph (AIG) format (Mishchenko et al., 2006), which adopts a Directed Acyclic Graph (DAG) structure with partially ordered nodes (Neto et al., 2019). Significant research efforts have aimed at embedding the inductive bias of DAGs into message-passing graph encoders (Thost and Chen, 2021) and graph transformers (Luo et al., 2024) to enhance model expressiveness. One contribution of our work extends these concepts by generating graph-level representations specifically optimized for AIGs.
@@
The current QoR prediction models often experience overfitting due to a limited number of publicly available circuits used during training, resulting in a lack of generalization, especially when tested on unseen circuits that differ in size and characteristics from the training circuits. Additionally, these models struggle with imbalanced input samples across different modalities; typically, there are more recipe available than AIGs. This disparity can lead to prominent overfitting, particularly due to the low expressiveness of graph encoders, which becomes exacerbated when the model excessively adapts to the encoder part handling recipe.
@@
Self-supervised learning (SSL) has recently emerged as a powerful framework to mitigate data scarcity and address expressiveness limitations in downstream tasks related to LSO (Chowdhury et al., 2023a; Wang et al., 2022). By pre-training circuit encoders and fine-tuning them for specific applications, SSL offers an efficient alternative to conventional data augmentation. Inspired by predictive SSL paradigms (Liu et al., 2022), this work introduces an auxiliary task that predicts intermediate QoRs, enabling a joint training approach that strengthens the supervisory signal and enhances model learning dynamics.
@@
Existing approaches face a critical limitation in the recipe encoder, which operates in isolation from AIG representations, leading to an inherent information bottleneck. This limitation becomes particularly evident when AIG and recipe embeddings are fused using conventional concatenation methods. This limitation motivated us to develop a decoder-only transformer model for training, employing a joint loss function. As a result, the model contextualizes the embeddings of heuristics using AIG embeddings through a Cross-Attention module, facilitating more effective integration of information.
@@ 2. Related Work
Several policy-based RL methodologies, such as those discussed in (Yang et al., 2022b; Chowdhury et al., 2024; Wu et al., 2022b), leverage a pre-trained function that correlates initial circuits and recipe with the ultimate QoR. This architecture is split into two primary branches: one for encoding AIGs and another for recipes, along with a fusion branch that maps the embeddings onto the predicted QoR. Similar efforts to forecast QoRs based on initial graphs and recipe are documented in (Chowdhury et al., 2021; Yang et al., 2022b, a; Wu et al., 2022a, b; Zheng et al., 2024), where various graph and sequence encoders are explored. Discussions abound concerning the most effective AIG encoders and methods for integrating sequence embeddings with graph embeddings. As a result, improvements in encoder and fusion modules are crucial for enhancing prediction performance. For example, (Chowdhury et al., 2021) utilizes a GCN to encode graphs and a CNN for recipes, employing both mean and max pooling for graph combination. Meanwhile, (Yang et al., 2022a) employs GraphSage for graph encoding and a transformer for recipe encoding. Similar to one of recent baselines, LOSTIN (Wu et al., 2022a), (Yang et al., 2022b) integrates a Graph Isomorphism Network (GINE) for graph learning with an LSTM for recipe encoding. Additionally, (Wu et al., 2022b) explores various graph encoders, such as PNA and pooling techniques, alongside virtual supernodes in conjunction with hierarchical learning and RL methods. In all aforementioned methods, recipe embeddings and Graph embeddings are fused using concatenation. The MLP decoder has been used to generate the predicted QoR after fusion. However, none of these models utilize an attention mechanism to fuse the two branches.
@@
$y_{k}$ The ground truth QoR of $k^{th}$ step
@@
Figure 1. Schematic overview of the LSOformer architecture during training and inference modes.
@@ 3.2.2. Final QoR prediction
(2) $loss=\sum_{k=1}^{M}MSE(\hat{y}_{k},y_{k})$
Where $y_{k}$ represents the ground truth QoR of $k^{th}$ step.
@@ 3.3.2. Level-wise Node Pooling
Table 2. Performance comparison of delay and area prediction models across datasets, evaluated using MAPE (Avg. ± Std.) (%).
@@ 3.3.4. Transformer Decoder
The mask $\bar{M}$ is typically an upper triangular matrix with zeros on and below the diagonal and $\infty$- ∞ above the diagonal for each sequence in the batch. This structure effectively imposes causality and ensures that each position in the sequence can only attend to itself and previous positions, not to any future positions.
@@
Where $\widetilde{h}^{i}=(\widetilde{u}_{1},\widetilde{u}_{2},...,\widetilde{u}_{M})$ is the output sequence of embeddings for input AIG and $i^{th}$ recipe; $\widetilde{u}_{i}\in\mathbb{R}^{2\times d_{h}}$ represents generated embedding correspoding to $i^{th}$ QoR.
Add & Norm: Each sub-layer (self-attention and feed-forward network) includes a residual connection followed by layer normalization:
@@
(14) $\hat{y}_{i}=MLP(\widetilde{h}^{i})$
Table 3. Statistics for benchmark datasets.
@@ 4.2. Datasets
Three datasets were employed for the experiments, each comprising a collection of circuit designs beside a set of recipe. Since each dataset is used in different studies, we chose not to combine them to simplify comparisons with baseline methods. Additionally, we report performance metrics individually for each dataset. Logic-level circuit designs in the datasets are referred to as Intellectual Property (IP). A consistent set of 1500 unique recipes borrowed from (Chowdhury et al., 2021) was applied across all datasets. Each recipe was constructed utilizing seven primary heuristics and their associated flags from the ABC toolkit. These heuristics include Balance, Rewrite (rw, rw -z), Refactor (rf, rf -z), and Re-substitution (rs, rs -z).
@@
Table 4. Ablation study on Recipe Encoder (Recipe Enc.) and AIG Encoder (AIG Enc.) using the Freeze Probing setup, evaluated with MAPE.
@@
Table 5. Ablation study on the decoder and SSL auxiliary task, evaluated with the MAPE metric. P.Q.T. stands for Predicting QoR trajectory.
@@ 4.3. Experimental Protocol
The efficacy of our model is assessed under both transductive and inductive settings to demonstrate its superiority over competing architectures. The dataset is split into training and validation sets with ratios of 0.66 and 0.33, respectively, for both setups. Additionally, zero-mean normalization is applied to normalize the final QoRs. Additionally, intermediate QoRs are normalized using the mean and standard deviation derived from the final QoRs. In the ablation study, we focus exclusively on the inductive setting, which presents the most challenging scenario for the LSO problem. In these experiments, given the challenges associated with predicting timing, delay are selected as the target variable for QoR prediction.
Inductive Setup (IP-Inductive): In this setup, the circuits used during the training phase are different from those employed in testing. It is noteworthy that all recipes from the sample set are observed and applied to the circuits introduced to the model during training.
Transductive Setup (Recipe-Inductive): In the transductive setup, each circuit is exposed during training, but separate recipes are used for testing. Specifically, the set of recipes is partitioned into test and validation subsets, and all Circuits are visible during both training and testing phases, albeit with different recipes.
@@ 4.4. Results
Table 6. Ablation study on different decoder layers for autoregressive and causal QoR trajectory prediction. Values represents MAPE metric.
@@ 4.4.1. Comparison Study Result
In Table 2, we evaluate the MAPE performance of various QoR prediction models against the baseline in multi trials across two setups for predicting delay and area. The baseline model is the architecture proposed in the OABCD paper (Chowdhury et al., 2021). In the IP-inductive setup, the LSOformer surpasses all existing models and outperforms the baseline by 5.74%, 4.35%, and 17.06% for the EPFL, OABCD, and PD datasets, respectively, in terms of delay prediction. For area prediction, LSOformer achieves a lower MAPE across all datasets compared to the baseline and nearly all models. However, on the OABCD dataset, the GNN-H model achieves the highest performance, demonstrating a 14% improvement over the baseline. This pattern holds in the recipe-inductive setup for both delay and area predictions across the three datasets, where LSOformer exceeds the baseline by 11.40%, 2.38%, and 12.82% for the EPFL, OABCD, and PD datasets in delay prediction, respectively.
Figure 2. Mean and circuit-wise performance comparison across test circuits for Area (a) and Delay (b).
@@ 4.4.2. Ablation Study
We assess the impact of the SSL auxiliary loss integrated to the training pipeline in IP-inductive setup using MAPE as evaluation metric. Similar to the benchmark SSL works (Liu et al., 2022), for the random setup, different parts of architecture are freezed to generate the lower bound of performance for SSL task. The difference between the results of random encoders and supervised models is considered the gap that the frozen model can potentially bridge in the best-case scenarios. Freezed LSOformer in Table 4, benefiting from QoR trajectory prediction, has performance near that of the supervised LSOformer and shows better results compared to the baseline, although it has not trained on final QoR.
@@
The effects of varying the embedding fusion and decoder configurations using different sequence decoders are explored in this work. The baseline model, which uses a shared MLP to predict all QoRs from initial to final, performs worse than alternative designs. Alternatively, separate MLPs are employed for each QoR, adopting a multi-task learning approach. This configuration significantly enhances performance across all datasets compared to the baseline. Additionally, we experimented with an auto-regressive LSTM module designed to train and infer models auto-regressively, thereby introducing causality to LSTM. This model demonstrates improvements over the baseline in all datasets except for OABCD. As can be seen in Table 6, the LSOformer surpasses both the baseline and all other methods across the three datasets, affirming the efficacy of our architectural design choices.